OpenStudy (anonymous):

The sum of three numbers is 137. The second number is four more than two times the first number. The Third number is five less than three times the first number. find the three numbers.

OpenStudy (anonymous):

x + (2x+4) +3x-5 = 137 x + 2x + 3x + 4-5 = 137 6x = 137-4+5 6x=138 x=138/6 x=23

OpenStudy (anonymous):

Therefore, 23, 50, 64